The cost of producing a certain item consist of P102 per unit for labor and material cost and P54 per  unit for other variable cost. The fixed cost per month amounts to P850,000. The item is sold at P740 each. 

a. Determine the break-even quantity per month. 

b. How many units must be produced each month in order that the net profit equals the cost?

c. What is the net profit it for a production of 4,000 units per month, in pesos? 




1
Expert's answer
2021-12-13T10:25:24-0500

Selling price (SP) = P740 per unit

Fixed cost (FC) = P850,000 per month

Variable cost (VC) = P102 + P54 = P156 per unit

a. Break-even quantity per month

"BEP=\frac{FC}{SP-VC}"


"BEP=\frac{850,000}{740-156}=1455"


b. Units where net profit equals the cost

"Total variable cost (TVC) = 156\times4000=P624,000"

"Total cost (TC) =TVC+FC=624,000+850,000=P1,474,000"


"Required sales=\frac{FC+Desiredprofit}{SP-VC}"


"Required sales=\frac{850000+1474000}{740-156}=3979"


This means 3,979 units need to be produced each month in order that the net profit equals the cost.


c. Net profit for a production of 4,000 units per month

"TC=P1,474,000"

"Total revenue (TR)=740\times4000=P2,960,000"

"Net profit (NP)= TR-TVC-FC"

"NP=2,960,000-624,000-850,000=P1,486,000"